### v1

@@ II.3 Equations of motion
where we introduce the polarization operator of all absorbers
@@
The last two terms are added to Eq. (8) following the input-output theory Courty and Reynaud (1992); Collett and Gardiner (1984), they describe the dissipation of the dipole momentum oscillations with the rate $\Gamma$; ${{\hat{b}}_{in}}$ is the Bose operator of the bath. The stationary solution of Eq. (8) is
@@
Using well-known properties of Dirac notations Lifshitz, Pitaevskii, and Berestetskii (1982) we come to commutation relations
@@
where ${{\hat{N}}_{0i}}={{\left|00\right\rangle}_{i}}{{\left\langle 00\right|}_{i}}$, ${{\hat{N}}_{1i}}={{\left|01\right\rangle}_{i}}{{\left\langle 01\right|}_{i}}$ are $i$-th absorber state population operators. Operators with different indexes $i$ commute. Using commutation relations (11), Hamiltonian (7) and polarisation (9) we write the equation of motion
@@
where ${{\hat{N}}_{\alpha}}=\sum\limits_{i=1}^{N}{{{{\hat{N}}}_{\alpha i}}}$, $\alpha=0,1$ are operators of populations of states of all absorbers. The last two terms in Eq. (12) describe the absorber polarization decay with the rate $\gamma$ and the noise with the Langevin force $\hat{F_{p}}$ of the absorber polarization baths.
Taking commutation relations $\left[{{{\hat{N}}}_{0i}},{{{\hat{\sigma}}}_{i}}\right]={{\hat{\sigma}}_{i}}$, Hamiltonian (7) and summing over all absorbers we obtain the equation of motion
@@
where the last two terms are due to the decay of the absorber excited state populations with the decay time $\tau$ and the Langevin force ${{\hat{F}}_{0}}$ corresponding to the population baths. The population operator of the upper states of absorbers is $\hat{N}_{1}=N_{m}-\hat{N}_{0}$.
III The stationary absorption
From here we use the semi-classical approach as, for example, in the laser theory Sargent, Scully, and Lamb (1974). So we neglect fluctuations, drop the bath operators, and replace operators with c-numbers (removing hats from notations). Calculating the semi-classical stationary solution of Eqs. (10) and (12) we find
@@ III The stationary absorption
Eq. (14) helps us to understand physical processes in the absorber-particle interaction. The stationary polarization of two-level absorbers in the resonant field without the particle is $p=-(\Omega_{m}/\gamma)a(N_{0}-N_{1})$. The particle near absorbers interacts with the external field with the rate $\Omega_{p}$, accepts the polarization, re-emits the field, and, by the re-emitted field, "transfers" the polarization (and the energy) to absorbers with the rate $\Omega_{int}$. Such polarization and the energy transfer is the plasmon-induced resonance energy transfer (PIRET) You, Ramakrishna, and Seideman (2018). PIRET is suppressed by the polarization decay: in the particle with the rate $\Gamma$ and in absorbers with the rate $\gamma$. This is why the PIRET particle-absorber polarization transfer rate is $\sim i\Omega_{int}\Omega_{p}/\Gamma\gamma$, as we see in Eq. (14). The multiplier $i$ in the last expression and in Eq. (14) is due to the well-known $\pi/2$ phase shift at the resonant interaction of oscillators Lifshitz and Landau (1982).
The absorber-particle interaction is a reciprocal process: the particle transfers the polarization to absorbers and vice-versa: absorbers transfer the polarization to the particle through the near field, causing the absorber polarization reduction. Such a reduction is described by the factor $1+\left({{N}_{0}}-{{N}_{1}}\right)/N_{int}$ in the denominator of Eq. (14).
In order to understand the physics and to characterize the stationary absorption we write, using Eq. (13), the energy conservation law: the rate $N_{1}/\tau$ of the energy relaxation from the upper states of absorbers is equal to the sum of the photon absorption rate from the external field $\sim\Omega_{m}$ and the energy exchange rates at the interaction of the particle and absorbers through plasmonic mode $\sim\Omega_{int}$
@@
The term $N_{1}$ in Eq. (17) describes the spontaneous emission, and the second term - a collective emission of absorbers to the plasmonic mode. These terms are similar to the cavity mode spontaneous emission and the collective emission terms in the laser theory André et al. (2019); Protsenko and Uskov (2022). We find from Eqs. (14), (15) and (17)
$\frac{{{N}_{1}}}{\tau}=$ ${\displaystyle\left(\frac{2\Omega_{m}^{2}}{\gamma}+\frac{2\Omega_{p}^{2}}{\Gamma{{N}_{int}}}\right)\frac{n({{N}_{0}}-{{N}_{1}})}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}}-}-$
$\frac{2\Omega_{int}^{2}}{\Gamma}\left({{N}_{1}}+{{\left|p\right|}^{2}}\right).$
where $n=|a|^{2}$ is the mean photon number in the volume $V$ of the external field. We express $n/V$ through the external field intensity later.
@@
and represent Eq. (III) in a compact form
@@
With the help of Eqs. (19) – (21) we identify processes affecting the absorber excited states population. The first process is the resonant absorption of the external field with the absorption coefficient $g_{m}$. The second process is the resonant absorption increase due to the nanoparticle near field. An external field energy is absorbed by the particle with the absorption coefficient $g_{p}$ and a part of this energy is transferred to absorbers by the near field with the transfer coefficient $1/N_{int}$. So the linear (i.e. independent on $N_{1}$) part of the absorption coefficient ${{\tilde{g}}_{m}}$ is $g_{m}+g_{p}/N_{int}$. The nonlinear term $\sim\tilde{\gamma}_{sp}$ in Eq. (21) is due to the collective emission of absorbers to plasmonic mode. The reason for the suppression factor $1+(N_{0}-N_{1})/N_{int}$ in the denominator of Eq. (20) is the transfer of the polarization and the energy from absorbers to the particle. The upper states population decays with the rate $\tilde{\gamma}_{sp}$ due to the spontaneous emission to the plasmonic mode, which is described by the last term on the right in Eq. (20).
IV Plasmon-enhanced absorption conditions and bistability
The resonant absorption in plasmonic mode is a nonlinear process. Besides the well-known resonant transition saturation, the nonlinearity is caused by the collective interaction of absorbers and the particle through the plasmonic mode leading to the energy exchange between absorbers and the particle. We will see how such nonlinear processes and the spontaneous emission to the plasmonic mode lead to the bistability in absorption and affect conditions for the best excitation of absorbers. Let us introduce some parameters.
IV.1 Parameters
@@
Here we use ${{{p}^{2}}}/{\hbar\Gamma}={{\alpha}_{p}}$, where ${{\alpha}_{p}}$ is the absolute value of the resonant polarisability of the metal nano-particle Protsenko et al. (2005). The external field intensity is $I=({c\eta}/{4\pi})\left\langle{{{\hat{\vec{E}}}}^{+}}\hat{\vec{E}}\right\rangle={c\hbar{{\omega}_{0}}n}/{V\eta}$. Here and below $c$ is the speed of light in vacuum. We take $I=\hbar{{\omega}_{0}}P$ then the external field intensity is $P={cn}/{V\eta}$ in the number of photons per the unit of area per second. So the number of photons in the external field per the unit of volume is ${n}/{V}={\eta P}/{c}$ and
@@
where ${{\lambda}_{0}}=2\pi/{{\omega}_{0}}$ is the absorbing field wavelength in vacuum, $\sigma_{p}$ is the particle resonant absorption cross-section. Similar way we find
@@
where $\alpha_{m}$ is the absolute value of the polarizability and $\sigma_{m}$ is the absorption cross-section of the two-level resonant absorber. It is convenient to represent the resonant absorber’s polarizability
@@
in terms of natural linewidth $\gamma_{sp}$ of the resonant transition Scully and Zubairy (1997); $k_{0}=\omega_{0}/c$ is the absorbed field wave-vector in vacuum. Using the parameters introduced above we express
@@
IV.2 Conditions for the absorption enhancement
We derive from the energy conservation law (20) and (21)
@@ IV.2 Conditions for the absorption enhancement
where ${{\sigma}_{eff}}({{N}_{1}})$ is a nonlinear (i.e. depended on $N_{1}$) cross-section of absorption
@@
In Eq. (26) we use $N_{0}+N_{1}=N_{m}$ and introduce the linear absorption cross-section enhancement factor
$\xi_{\sigma}=\frac{1+{\alpha_{p}^{2}}/{4{{r}^{6}}}}{1+\gamma\tau\left({\alpha_{p}\alpha_{m}}/{2{{r}^{6}}}\right)}$ (27)
The nominator of (27) is greater than 1 because of the direct energy transfer from the nanoparticle to absorbing molecules. Such transfer is described by the factor $\alpha_{p}^{2}/4r^{6}$. The denominator of Eq. (27) is greater than 1 because of the spontaneous emission to the LPR mode. Such spontaneous emission reduces the absorption cross-section. $\sigma_{eff}$ is reduced relatively to $\xi_{\sigma}\sigma_{m}$ because of the collective spontaneous emission of absorbers to the LPR mode and the energy transfer back from emitters to the particle. Both these processes make the denominator in Eq. (26) greater than 1.
Conditions, necessary for the absorption enhancement by the LPR follow from Eqs (26), (27). The absorption enhancement is, when $\sigma_{eff}(N_{1})>\sigma_{m}$. Suppose, in the beginning, that the denominator in Eq. (26) is close to 1, so the collective emission to the LPR mode and the energy transfer from absorbers to the nanoparticle are relatively small. This is true at the saturation of the absorbing transitions, when $N_{1}$ is close to $N_{m}/2$, or at a small number $N_{m}$ of absorbers, when
@@
When Eq. (28) is true, the absorption cross-section is enhanced at $\xi_{\sigma}>1$ which is when
@@
or, if we insert the expression (24) for $\alpha_{m}$ to (29), when $\alpha_{p}>\alpha_{p}^{min}=3\gamma_{sp}\tau/2\eta k_{0}^{3}$. The relation (29) is a necessary condition for the plasmon-stimulated absorption enhancement (found in the semiclassical approach). We note that $2\gamma\tau\geq 1$, so it must be $\alpha_{p}>\alpha_{m}$ for the absorption enhancement. So absorption enhancement is possible when the nanoparticle absorbs the field better than an emitter, which has been mentioned, for example, in Cortés (2017).
However the nanoparticle absorption can not be too large. Otherwise, the absorption by molecules will be suppressed due to their collective emission to the LPR mode and the energy transfer from molecules to the particle. So combining conditions (28) and (29) we obtain the interval of $\alpha_{p}$, where the the absorption enhancement by LPR is possible
@@
It is interesting to note that the fact of the absorption enhancement does not depend on the distance $r$ between the absorbers and the particle (as far as absorbers are in the nanoparticle near the field zone). However, the magnitude of the enhancement depends on $r$, as we will see.
IV.3 Bistability in the stationary absorption
Suppose the number $N_{m}$ of absorbing molecules is so large that $N_{m}{\alpha_{p}\alpha_{m}}/{2{{r}^{6}}}>1$. Then the absorption is strongly nonlinear, because of the collective effects, the energy transfer between absorbers and the particle, which makes a significant effect of $N_{1}$-dependent denominator in Eq. (26). The denominator is decreased with $N_{1}$, so the absorption is increased with $N_{1}$ at certain parameter values. As a result the low and the high absorption regimes are possible, they overlap at certain parameters making the bistability in absorption.
In order to investigate the bistability we introduce dimensionless absober-nanoparticle interaction $A={\alpha_{p}\alpha_{m}}{N}_{m}/{4{{r}^{6}}}$, the external field intensity ${{P}_{eff}}=\xi_{\sigma}\sigma_{m}P\tau$ parameters, and consider the upper state population per absorber $n_{1}={N_{1}}/{N_{m}}$. With these parameters we write a cubic equation on $n_{1}$
@@ IV.3 Bistability in the stationary absorption
instead of Eq. (25). Proceeding a standard analysis of the catastrophe theory Poston and Stewart (1996) we find that Eq. (31) has three solutions if the number $N_{m}$ of absorbers is so large that
@@
Note that the condition (32) is opposite to the small collective emission and the small absorber to particle energy transfer condition (28). Thus the collective emission and the absorber to particle energy transfer must be sufficiently large for the bistability.
The bistability domain is restricted by the $P_{\pm}(A)$ curves in $A$, $P_{eff}$ parameter space
@@
Figure 2: The absorption bistability is for values of $A$ and $P_{eff}$ between the curves $P_{\pm}(A)$.
V Examples of absorption enhancement and the bistability

### v2

@@ 2.3 Equations of motion
where we introduce the polarisation operator of all absorbers
@@
The last two terms are added to Eq. (8) according to the input-output theory [49, 50], they describe the dissipation of the dipole momentum oscillations with the rate $\Gamma$; ${{\hat{b}}_{in}}$ is the Bose operator of the bath. The stationary solution of Eq. (8) is
@@
Using well-known properties of Dirac notations [46], we arrive at the commutation relations
@@
where ${{\hat{N}}_{0i}}={{\left|00\right\rangle}_{i}}{{\left\langle 00\right|}_{i}}$, ${{\hat{N}}_{1i}}={{\left|01\right\rangle}_{i}}{{\left\langle 01\right|}_{i}}$ are ith absorber state population operators. Operators with different indices $i$ are commute. Using the commutation relations (11), Hamiltonian (7), and polarisation (9) we write the equation of motion
@@
where ${{\hat{N}}_{\alpha}}=\sum\limits_{i=1}^{N}{{{{\hat{N}}}_{\alpha i}}}$, $\alpha=0,1$ are operators of populations of states of all absorbers. The last two terms in Eq. (12) describe the absorber polarisation decay with the rate $\gamma$ and the noise with the Langevin force $\hat{F_{p}}$ of the absorber polarisation baths.
Taking the commutation relations $\left[{{{\hat{N}}}_{0i}},{{{\hat{\sigma}}}_{i}}\right]={{\hat{\sigma}}_{i}}$, Hamiltonian (7), and summing over all absorbers we obtain the equation of motion
@@
where the last two terms are due to the decay of the absorber excited state populations with the decay time $\tau$ and the Langevin force ${{\hat{F}}_{0}}$ corresponding to the population bath. The population operator of the absorber upper states is $\hat{N}_{1}=N_{m}-\hat{N}_{0}$.
3 The stationary absorption
From here we use the semi-classical approach as, for example, in the laser theory [31]. We neglect fluctuations, drop the bath operators, and replace the operators with c-numbers (removing the hats from the notations). Calculating the semi-classical stationary solution of Eqs. (10) and (12) we find
@@ 3 The stationary absorption
Eq. (14) helps us to understand the physical processes in the absorber-particle interaction. The stationary polarisation of two-level absorbers in the resonant field without the particle is $p=-(\Omega_{m}/\gamma)a(N_{0}-N_{1})$. The particle near the absorbers interacts with the external field at the rate $\Omega_{p}$, accepts the polarisation, re-emits the field, and through the re-emitted field ”transfers” the polarisation (and the energy) to absorbers with the rate $\Omega_{int}$. Such polarisation and the energy transfer is called plasmon induced resonance energy transfer (PIRET) [24]. PIRET is suppressed by the polarisation decay: in the particle at the rate $\Gamma$ and in the absorbers at the rate $\gamma$. Therefore, the PIRET particle-absorber polarisation transfer rate is $\sim i\Omega_{int}\Omega_{p}/\Gamma\gamma$, as we see in Eq. (14). The multiplier $i$ in the last expression and in Eq. (14) is due to the well-known $\pi/2$ phase shift in the resonant interaction of oscillators [51].
The interaction between the absorber and the particle is a two-way process: the particle transfers the polarisation to the absorber and vice versa: the absorber transfers the polarisation to the particle through the near field, causing a reduction in the polarisation of the absorber. Such a reduction is described by the factor $1+\left({{N}_{0}}-{{N}_{1}}\right)/N_{int}$ in the denominator of Eq. (14).
To understand the physics and to characterise the stationary absorption we write the energy conservation law using Eq. (13): the rate $N_{1}/\tau$ of the energy relaxation from the upper states of the absorbers is equal to the sum of the photon absorption rate from the external field $\sim\Omega_{m}$ and the energy exchange rates at the interaction of the particle with the absorbers through the plasmonic mode $\sim\Omega_{int}$
@@
The term $N_{1}$ in Eq. (17) describes the spontaneous emission, and the second term describes a collective emission of absorbers to the plasmonic mode. These terms are similar to the cavity mode spontaneous emission and the collective emission terms in the laser theory [28, 29]. From Eqs. (14), (15) and (17) we find
$\frac{{{N}_{1}}}{\tau}=\displaystyle\left(\frac{2\Omega_{m}^{2}}{\gamma}+\frac{2\Omega_{p}^{2}}{\Gamma{{N}_{int}}}\right)\frac{n({{N}_{0}}-{{N}_{1}})}{1+\left({{N}_{0}}-{{N}_{1}}\right)/{{N}_{int}}}-\frac{2\Omega_{int}^{2}}{\Gamma}\left({{N}_{1}}+{{\left|p\right|}^{2}}\right)$ (18)
where $n=|a|^{2}$ is the mean number of photons in the volume $V$ of the external field. We will later express $n/V$ in terms of the external field intensity.
@@
and represent Eq. (18) in a compact form
@@
Using Eqs. (19) – (21) we identify processes that affect the absorber excited state population. The first process is the resonant absorption of the external field with the absorption coefficient $g_{m}$. The second process is the resonant absorption enhancement due to the nanoparticle near field. An external field energy is absorbed by the particle with the absorption coefficient $g_{p}$ and a part of this energy is transferred to the absorbers by the near field with the transfer coefficient $1/N_{int}$. Thus the linear (i.e. independent of $N_{1}$) part of the absorption coefficient ${{\tilde{g}}_{m}}$ is $g_{m}+g_{p}/N_{int}$. The non-linear term $\sim\tilde{\gamma}_{sp}$ in Eq. (21) is due to the collective emission of absorbers to plasmonic mode. The reason for the suppression factor $1+(N_{0}-N_{1})/N_{int}$ in the denominator of Eq. (20) is due to the transfer of the polarisation and the energy from the absorbers to the particle. The upper state population decays at the rate $\tilde{\gamma}_{sp}$ due to the spontaneous emission to the plasmonic mode, which is described by the last term on the right in Eq. (20).
4 Plasmon-enhanced absorption conditions and bistability
Resonant absorption in the plasmonic mode is a non-linear process. In addition to the well-known resonant transition saturation, the nonlinearity is caused by by the collective interaction of the absorbers and the particle through the plasmonic mode, leading to an energy exchange between the absorbers and the particle. We will see how such nonlinear processes and the spontaneous emission to the plasmonic mode lead to the bistability of the absorption and affect the conditions for the best excitation of the absorbers. Let us introduce some parameters.
4.1 Parameters
@@
Here we use ${{{p}^{2}}}/{\hbar\Gamma}={{\alpha}_{p}}$, where ${{\alpha}_{p}}$ is the absolute value of the resonant polarisability of the metal nanoparticle [25]. The external field intensity is $I=({c\eta}/{4\pi})\left\langle{{{\hat{\vec{E}}}}^{+}}\hat{\vec{E}}\right\rangle={c\hbar{{\omega}_{0}}n}/{V\eta}$. Here and below $c$ is the speed of light in vacuum. If we take $I=\hbar{{\omega}_{0}}P$ then the external field intensity is $P={cn}/{V\eta}$ in the number of photons per the unit of area per second. So the number of photons in the external field per the unit of volume is ${n}/{V}={\eta P}/{c}$ and
@@
where ${{\lambda}_{0}}=2\pi/{{\omega}_{0}}$ is the wavelength of the absorbing field in vacuum, $\sigma_{p}$ is the absorption cross section of the particle resonance. Similarly, we find
@@
where $\alpha_{m}$ is the absolute value of the polarisability and $\sigma_{m}$ is the absorption cross section of the two-level resonant absorber. It is convenient to express the polarisability of the resonant absorber
@@
in terms of the natural linewidth $\gamma_{sp}$ of the resonant transition [52]; $k_{0}=\omega_{0}/c$ is the absorbed field wave vector in vacuum. Using the parameters introduced above, we express
@@
4.2 Conditions for the absorption enhancement
From the energy conservation law (20) and (21) we derive
@@ 4.2 Conditions for the absorption enhancement
where ${{\sigma}_{eff}}({{N}_{1}})$ is a non-linear (i.e. dependent on $N_{1}$) absorption cross section
@@
In Eq. (26) we use $N_{0}+N_{1}=N_{m}$ and introduce the linear absorption cross section enhancement factor
$\xi_{\sigma}=\frac{1+{\alpha_{p}^{2}}/{4{{r}^{6}}}}{1+\gamma\tau\left({\alpha_{p}\alpha_{m}}/{2{{r}^{6}}}\right)}.$ (27)
The nominator of (27) is greater than 1 because of the direct energy transfer from the nanoparticle to the absorbing molecules. Such transfer is described by the factor $\alpha_{p}^{2}/4r^{6}$. The denominator of Eq. (27) is greater than 1 because of the spontaneous emission to the LPR mode. Such spontaneous emission reduces the absorption cross section. $\sigma_{eff}$ is reduced relative to $\xi_{\sigma}\sigma_{m}$ due to the collective spontaneous emission of absorbers to the LPR mode and the energy transfer back from absorbers to the particle. These two processes make the denominator in Eq. (26) greater than 1.
The conditions, necessary for the absorption enhancement by the LPR follow from Eqs (26), (27). The absorption enhancement is present when $\sigma_{eff}(N_{1})>\sigma_{m}$. Assume at the beginning, that the denominator in Eq. (26) is close to 1, so that the collective emission to the LPR mode and the energy transfer from the absorbers to the nanoparticle are relatively small. This is the case at the saturation of the absorbing transitions, when $N_{1}$ is close to $N_{m}/2$, or at a small number of $N_{m}$ of absorbers, when
@@
If Eq. (28) holds, the absorption cross section is enhanced at $\xi_{\sigma}>1$ which is when
@@
or, if we insert the expression (24) for $\alpha_{m}$ into (29), when $\alpha_{p}>\alpha_{p}^{min}=3\gamma_{sp}\tau/2\eta k_{0}^{3}$. The relation (29) is a necessary condition for the plasmon-stimulated absorption enhancement (found in the semiclassical approach). Note that $2\gamma\tau\geq 1$, so $\alpha_{p}>\alpha_{m}$ is required for absorption enhancement. So absorption enhancement is possible if the nanoparticle absorbs the field better than an absorber, as mentioned for example in [11].
However, the absorption of the nanoparticles must not be too high. Otherwise, the absorption by the molecules will be suppressed due to their collective emission to the LPR mode and the energy transfer from the molecules to the particle. Thus, by combining conditions (28) and (29) we obtain the interval of $\alpha_{p}$, where the absorption enhancement by LPR is possible
@@
It is interesting to note that the fact of the absorption enhancement does not depend on the distance $r$ between the absorbers and the particle (as long as the absorbers are in the nanoparticle near the field zone). However, the magnitude of the enhancement does depend on $r$, as we will see.
4.3 Bistability in the stationary absorption
Suppose the number $N_{m}$ of absorbing molecules is such that $N_{m}{\alpha_{p}\alpha_{m}}/{2{{r}^{6}}}>1$. Then the absorption is strongly non-linear due to the collective effects, the energy transfer between the absorbers and the particle, which has a significant effect of the $N_{1}$-dependent denominator in Eq. (26). The denominator decreases with $N_{1}$, so the absorption increases with $N_{1}$ for certain parameter values. As a result, the low and the high absorption regimes are possible, overlapping at certain parameters, creating the bistability in the absorption.
To study the bistability we introduce the dimensionless nanoparticle-absorber interaction $A={\alpha_{p}\alpha_{m}}{N}_{m}/{4{{r}^{6}}}$, the external field intensity ${{P}_{eff}}=\xi_{\sigma}\sigma_{m}P\tau$ parameters, and consider the upper state population per absorber $n_{1}={N_{1}}/{N_{m}}$. With these parameters we write a cubic equation for $n_{1}$
@@ 4.3 Bistability in the stationary absorption
instead of Eq. (25). Following a standard analysis of the catastrophe theory [53], we find that Eq. (31) has three solutions when the number $N_{m}$ of absorbers is large enough that
@@
Note that the condition (32) is opposite to the small collective emission and the small absorber-to-particle energy transfer condition (28). Thus the collective emission and the absorber to particle energy transfer must be sufficiently large for the bistability.
The bistability domain is bounded by the $P_{\pm}(A)$ curves in $A$, $P_{eff}$ parameter space
@@
Figure 2: Absorption bistability for values of $A$ and $P_{eff}$ lie between the curves $P_{\pm}(A)$.
5 Examples of absorption enhancement and the bistability